OpenStudy (anonymous):

give an indirect proof that states that the sqrt(2400) is not equal to 49

OpenStudy (unklerhaukus):

\[49^2=49\times 49=2401≠2400\]

OpenStudy (unklerhaukus):

\[ \therefore \sqrt{2400}≠49\]